OpenStudy (adamaero):

MEDAL even for HELP - If 1,000 units are required per year, which is least costly? Alternative Resources Required A Skilled Operator at $25 per hour; 35 minutes per unit of skilled labor time B Unskilled operator at $12 per hour plus automated equipment costing $50,000; 10 minutes of operator time per unit

OpenStudy (greencat):

Where'd you get 35 from?

OpenStudy (adamaero):

min/unit

OpenStudy (greencat):

I don't quite get what you mean. @satellite73

OpenStudy (greencat):

@HelpBlahBlahBlah

OpenStudy (greencat):

@mathstudent55

OpenStudy (adamaero):

Alternative B at $12/unit Alternative A at $10.50/unit I tried 35*1000 = 35000 35000/60 = num num*25 = 14.583 (or 14.6 per unit) $14.6/unit is wrong What am I doing wrong?

OpenStudy (adamaero):

$2000+$50,000 for alternative B?

OpenStudy (adamaero):

but that doesn't make sense at $52/unit

OpenStudy (greencat):

@aaronq